OpenStudy (anonymous):

(x+14.5)^2/.25 - (y+9.5)^2/30.25 =1 convert to y=. do not simplify

OpenStudy (gw2011):

In order to find y some simplification has to occur. Here is the solution I came up with: Multiply both sides of the equation by 30.25 in order to eliminate the denominator. Therefore, you get: 121(x+14.5)^2 - (y+9.5)^2 = 30.25 (y+9.5)^2 = 121(x+14.5)^2 - 30.25 Getting the square root you get : y+9.5 = 11(x+14.5) - 5.5 y = 11(x+14.5) - 5.5 - 9.5 y = 11(x+14.5) - 15 y = 11x + 159.5 - 15 y = 11x + 144.5 You can follow the steps above and see at what step you want to stop further simplication. Since the problem states to convert to y, then the above steps show that process.